OpenStudy (anonymous):

solve. you invest $2000 in a bank account for 5 yrs at an annual interest rate of 5%. how much will your account be worth after 5yrs if it is compunded four times a year, to the nearest $0.01? how much more will your account be worth after five years if it is compunded continuously?

OpenStudy (anonymous):

just multiply it by 0.05 for each year.

OpenStudy (anonymous):

2000 times 0.05 = x

OpenStudy (anonymous):

x times 0.05 = y ... etc until you get to the five years.

OpenStudy (anonymous):

1.05, my bad.

OpenStudy (anonymous):

2000+1.05x=y for part 1

OpenStudy (anonymous):

For the first question: A=$2000 (1+(.05)/4)^4=$2000 (1.0125)^4=$2102.89 For the second question: A=$2000e^(rate*time)=$2000e^(.05*5)=$2568.05

OpenStudy (anonymous):

Cristiano and Johnny are both painfully mistaken.

OpenStudy (anonymous):

Lol, you're right. I've not done this in a while.